SOLVE dx/dy=1/x +1/y;


 

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
\frac{1}{y'} = \frac{1}{x} + \frac{1}{y}
It is a non- linear 1storder differential equation & its not in the syllabus of JEE. We have linear 1storder differential equation in JEE.
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty